# "I’m not for sale" – Perceptions and limited awareness of privacy risks by digital natives about location data

### v1

@@
Abstract
Although mobile devices benefit users in their daily lives in numerous ways, they also raise several privacy concerns. For instance, they can reveal sensitive information that can be inferred from location data. This location data is shared through service providers as well as mobile applications. Understanding how and with whom users share their location data – as well as users’ perception of the underlying privacy risks –, are important notions to grasp in order to design usable privacy-enhancing technologies. In this work, we perform a quantitative and qualitative analysis of smartphone users’ awareness, perception and self-reported behavior towards location data-sharing through a survey of $n=99$ young adult participants (i.e., digital natives). We compare stated practices with actual behaviors to better understand their mental models, and survey participants’ understanding of privacy risks before and after the inspection of location traces and the information that can be inferred therefrom.
@@ 1 Introduction
Smartphones have become the most popular electronic devices, used by 95% of all internet users [^1] https://datareportal.com/global-digital-overview. The widespread adoption of these mobile devices with geolocation capabilities makes users’ location traces [^2] That is, location data collected by Google from smartphones which gives more accurate and reliable information than other kinds of location data collection widely accessible to mobile services. Location is the most collected personal data by mobile applications (Achara et al. 2013), and while the exploitation of this data presents an obvious benefit to users in their daily lives – through personalized services, and potentially to the society with crowdsourcing initiatives (Stevens and D’Hondt 2010) –, an uncontrolled exploitation of this data continue to raise privacy concerns.
@@
This location data can indeed reveal a lot of sensitive information about users and represents unique privacy risks and implications compared to other types of data shared via smartphones. For example, it has been shown that it is possible to infer personality traits, religious affiliation, sexual orientation, or health status (Gambs, Killijian, and del Prado Cortez 2010). Additionally, the uniqueness of location data provides a sort of fingerprint specific to each individual that can be used to re-identify users in anonymized data, i.e., little a priori knowledge about a user can be exploited to discriminate them in a large set of anonymous location data. For instance, journalists were able to re-identify and track the whereabouts of the former US president Trump from a large dataset (Stuart A. and Charlie 2019), and anonymous data was used to pinpoint Muslim cab drivers (Lorenzo 2015). The attack surface of the exploitation of personal data is not fully discovered. Regularly, new threats appear and are maliciously exploited for influential campaigns and interfering in political election [^3] https://en.wikipedia.org/wiki/Troll˙farm, or discovering sensitive information [^4] https://www.nytimes.com/2018/01/29/world/middleeast/strava-heat-map.html.
@@
With this increasing exposure of privacy risks, understanding how young users share their data, and the extent to which they are aware of security and privacy risks, are important properties to assess these risks and to develop effective Privacy and Transparency Enhancing Technologies matching current users’ expectations to increase adoption. Although user perceptions of technology (Beckwith 2003) and the privacy paradox [^5] The privacy paradox refers to self-reported concerns about privacy appear to be in contradiction with often careless online behaviors. (Barth et al. 2019; Kang and Jung 2021) have received a lot of attention, users’ self-reported behaviors in mobility contexts associated with smartphones, and the impact of an awareness demonstration platform have been less studied in the academic literature. To fill this gap, this work addresses two research questions:
  • RQ1
@@
    What are the perceptions and the understanding of young users’ privacy and its protection in a mobility context?
  • RQ2
@@
    What is the impact of a demonstrator for the visualization of location traces and associated privacy risks on these perceptions and understanding?
@@ 2.1 Privacy and location
The privacy issues raised by location data gained a lot of traction in the last decade (Primault et al. 2019). In particular, user location traces extracted from various data have been shown to be highly unique (De Montjoye et al. 2013; Zang and Bolot 2011). This high uniqueness may act as a digital fingerprint and lead to the re-identification of users if their traces are associated with external knowledge. This uniqueness does not only concern location traces but characterizes all traces generated by a human  (Andreas et al. 2016; Eckersley 2010; Rebekah and Rachel 2016), and can generate a risk of re-identification. Several cases of re-identification have been documented, for instance the re-identification of individuals from web search queries (Barbaro and Zeller 2006), taxi logs (Zhang and Wang 2016), or the notorious case of Governor William Weld using medical information (Barth-Jones 2012).
@@
The uniqueness and risk of re-identification is not the unique threat related to location traces. Recent works have demonstrated that location is a very rich contextual information and leads to a strong inferential potential in terms of information that can be learned about individuals (Boutet and Gambs 2019). For instance, location traces can reveal Points Of Interest (POI) of users such as their home and workplaces (Gambs, Killijian, and del Prado Cortez 2010), their race and gender (Zhong et al. 2015), their social network (Sharad and Danezis 2014), it can be used to predict their location patterns (Sadilek and Krumm 2012), to link accounts of the same user across different datasets (Riederer et al. 2016a), and infer even more sensitive information such as their religion or personality traits [^6] Note that the processing of such sensitive information is prohibited by Article 9(1) of the General Data Protection Regulation (GDPR) on the processing of special categories of personal data..
@@
Due to the large adoption of mobile devices, the location data of users is extensively collected and shared (Almuhimedi et al. 2015; Englehardt and Narayanan 2016). The uncontrolled usage of this information can have an important impact on users such as unfair price discrimination (Mikians et al. 2012). The increasing use of connected devices collecting personal data and the lack of transparency on how the data is actually exploited raise privacy concerns. Only a handful of tools have been proposed to improve user awareness about the potential risk of revealing their location. For instance, (Riederer et al. 2016b) propose a tool to inspect the potential of location data, while (Boutet and Gambs 2019) show the impact of protection mechanisms of the inference capabilities using a demonstration platform.
2.2 Perceptions of privacy and privacy controls
@@
Giving users the benefits of location services on their mobile devices while preserving their privacy is an ongoing challenge, as evidenced by mobile OSs iterating on the user interfaces for location notices and control every few years. For example, iOS allows users to select whether apps get fine- or coarse-grained location data [^7] https://support.apple.com/guide/iphone/control-the-location-information-you-share-iph3dd5f9be/ios, and will present pop-ups when apps continually access location services in the background. Similarly, Android apps now have to request background location access separately from general location usage [^8] https://developer.android.com/about/versions/11/privacy/location, and the OS will automatically revoke unused permissions from apps [^9] https://developer.android.com/about/versions/11/privacy/permissions. Even though mobile operating systems regularly improve user interfaces, the opaque privacy controls of location services still face criticism [^10] https://www.attorneygeneral.gov/taking-action/attorney-general-josh-shapiro-announces-391-million-settlement-with-google-over-location-tracking-practices/. (Balash et al. 2022) explore the users’ perceptions regarding access to Google accounts by mobile applications and formulate design recommendations to improve the current third-party management tools offered by Google, such as tracking recent access, automatically revoking access due to app disuse, and providing permission controls. (Wijesekera et al. 2015) analyze authorization preferences in different usage contexts and suggest determining the situations in which users would like to be confronted with security decisions.
@@
Users’ perceptions of the risks and benefits of technologies can determine their willingness to adopt them (Poikela and Kaiser 2016). More specifically, people are more likely to accept potentially invasive technology if they think its benefits will outweigh its potential risks (Beckwith 2003). Due to the massive adoption of location-enabled mobile applications, this fact suggests that users’ perception of this trade-off is more in favor of the benefits than the potential risks. This attitude however depends on the perception of the said privacy risks. Studies have shown that users of mobile phones are often unaware of the data collected by apps running on their devices, and that a majority of users restrict some of their permissions (Almuhimedi et al. 2015) following a better awareness of data collection. Other studies explored the privacy paradox (Barth et al. 2019; Kang and Jung 2021) where self-reported concerns about privacy appear to be in contradiction with often careless online behaviors. However, another study focuses on university community (Gamarra et al. 2019) and show that this population of users does not have a genuine concern regarding the privacy of their geolocation data. Note that this paper did not study location data in mobility contexts associated with smartphones as we do.
@@
Analyzing behavior and understanding users’ perceptions are also important notions to grasp in order to further design Privacy and Transparency Enhancing Technologies (PETs and TETs). For instance, (Kaushik et al. 2021) conducted an online survey to understand people’s perspectives on solely automated decision-making. They then formulate recommendations on how to design such systems. In a similar line of work, (Islami, Fischer-Hübner, and Papadimitratos 2022) conducted in-depth semi-structured interviews with 17 Swedish drivers to analyse their privacy perceptions and preferences for intelligent transportation systems, then to provide recommendations for suitable predefined privacy options. Several other works (Zufferey et al. 2023; Velykoivanenko et al. 2021; Mink et al. 2022) studied the perceptions of privacy and utility of users related to fitness-trackers and demonstrated a high potential for data minimization (i.e., reducing the volume of data sent to service provider). (Debatin et al. 2009), in turn, investigated Facebook users’ awareness of privacy issues and perceived benefits and risks of utilizing Facebook and recommended better privacy protection, higher transparency and more education about the risks of posting personal information to reduce risky behavior. (Bielova et al. 2024) analysed the behaviors of websites’ visitors through a study of the impact of dark patterns on consent decisions. (Veys et al. 2021) explored whether current data downloads (such as Google Takeout) actually achieve the transparency goals embodied by the right of access. Most participants indicated that current offerings need improvement to be useful, emphasizing the need for better filtration, visualization, and summarizing to help them hone in on key information. However, none of these related work specifically targets users’ behavior and perception about data-sharing of location data. Although (Martin and Nissenbaum 2019) clearly addressed users’ perceptions of location data, it is to be noted that it does so from an information science standpoint, closer to sociology and not from a UX/usability one. It nonetheless offers a relevant account for the interested reader.
@@
Related to location data, (Farke et al. 2021) specifically analysed the user perceptions and reactions to Google’s My Activity and how this web history dashboard increases or decreases end-users’ concerns and benefits regarding data collection. Their results show that participants were surprised by the volume and detail of the collected data, but most of them were significantly more likely to be both 1) less concerned about data collection and 2) to view data collection more beneficially. However, this dashboard does not present any risks such as possible sensitive inferences associated with location traces or places visited. By also presenting the risks, our study shows that users are more concerned about privacy issues.
3 Methodology
@@
In a nutshell, the first questionnaire (hereinafter “behavior questionnaire”) aimed at assessing their self-reported practices, perceptions, and pre-conceived ideas about privacy. Following this first questionnaire, participants were then invited to use a demonstrator (denoted “demonstration” in what follows) which analyses location traces and reports several inferences as well as the impact of a protection mechanism using differential privacy (i.e., geo-indistiguishability (André s et al. 2013)). The second questionnaire (hereinafter “perception questionnaire”) addressed their assessment of the risks.
3.1 Recruitment
@@ 3.2 Ethical considerations
Second, we elicited answers from participants, which is also personal data. We took as many precautions as needed to guarantee 1) informed consent (see Section B.1 for the consent form, and as noted in Section 3.1 they had the possibility to decline without any penalty), and 2) security of data storage and processing (authors only communicated using end-to-end encrypted tools, and the survey was conducted on an EU-based online tool stamped GDPR-compliant). One participant refused to answer the questionnaires.
@@
3.3 Design of the Questionnaires
We shortly describe in this section the organization of the two questionnaires. The interested reader will find their full content in Appendix B. Each section is succinctly described in what follows:
@@ 3.3 Design of the Questionnaires
3.4 Risks analysis - demonstration
After the first behavior questionnaire, participants had to explore a demonstration platform informing them about privacy risks linked to location traces through example data. We use the Boutet and Gambs (2019) demonstration platform. Participants were asked to use the platform to explore these location traces and to inspect the information that can be inferred. The demonstration typically offers to visualize one’s traces per day (see Figure 1), which can raise awareness about how easy it is to identify Points of Interests (POIs) – such as the home, the workplace or any other visited places during the day –, and presents metadata (e.g., address, category of the place, attendance statistics).
Figure 1: Screenshot of the demonstration’s interface. Clicking on a POI yields attendance statistics.
The demonstration also offers a defense mechanism based on Differential Privacy (Dwork 2006) in the form of a slider, which controls the level of noise injected into the location data (see Figure 2). This feature is presented as a means to stimulate interest in privacy protection, and to raise awareness of the degradation of data quality when we increase protection with greater noise (to showcase the utility and privacy trade-off).
Figure 2: Illustration of the defense mechanism. Here the red lines show the initial traces, and the black lines represent the traces after the application of the noise.
Finally, the demonstration highlights the risks raised by analyzing location traces and inferring gender, big five personality traits, home, and workplace (see Figure 3). Given the large number of inferences of sensitive information from location data reported in the literature (e.g., people encountered, religious affiliation, sexual orientation, or health status (Gambs, Killijian, and del Prado Cortez 2010)), only the presentation of these inferences fits fairly well into a representative risk of targeting (i.e., based on predictions computed on data collected and potentially exchanged between different parties).
Figure 3: Inference performed by the platform. The left-hand side shows inference from raw data while the right-hand side presents the reduced risks after the application of the defense mechanism (i.e., geo-indistinguishability (André s et al. 2013)). The noisy version hardly yields the gender, smooths the personality traits, makes the daily and weekly attendance statistics unusable, and infers inaccurate POIs.
3.5 Data analysis
We performed a thematic analysis on the free text answered in the study (see Tables 1 and 2 in Appendix C) to analyse qualitative data. We took inspiration from the steps described in (Braun and Clarke 2006):
  1. 1.

### v3

@@
Abstract.
Although mobile devices benefit users in their daily lives in numerous ways, they also raise several privacy concerns. For instance, they can reveal sensitive information that can be inferred from location data. This location data is shared through service providers as well as mobile applications. Understanding how and with whom users share their location data as well as users’ perception of the underlying privacy risks, are important notions to grasp in order to design usable privacy-enhancing technologies. In this work, we perform a quantitative and qualitative analysis of smartphone users’ awareness, perception and self-reported behavior towards location data-sharing through a survey of $n=99$ young adult participants (i.e., digital natives). We compare stated practices with actual behaviors to better understand their mental models, and survey participants’ understanding of privacy risks before and after the inspection of location traces and the information that can be inferred therefrom.
@@ 1. Introduction
Smartphones have become the most popular electronic devices, used by 95% of all internet users [^1] https://datareportal.com/global-digital-overview. The widespread adoption of these mobile devices with geolocation capabilities makes users’ location traces [^2] That is, location data collected by Google from smartphones which gives more accurate and reliable information than other kinds of location data collection widely accessible to mobile services. Location is the most collected personal data by mobile applications (Achara et al., 2013), and while the exploitation of this data presents an obvious benefit to users in their daily lives – through personalized services, and potentially to the society with crowdsourcing initiatives (Stevens and D’Hondt, 2010) –, an uncontrolled exploitation of this data continue to raise privacy concerns.
@@
This location data can indeed reveal a lot of sensitive information about users and represents unique privacy risks and implications compared to other types of data shared via smartphones. For example, it has been shown that it is possible to infer personality traits, religious affiliation, sexual orientation, or health status (Gambs, Killijian, and del Prado Cortez, 2010). Additionally, the uniqueness of location data provides a sort of fingerprint specific to each individual that can be used to re-identify users in anonymized data, i.e., little a priori knowledge about a user can be exploited to discriminate them in a large set of anonymous location data. For instance, journalists were able to re-identify and track the whereabouts of the former US president Trump from a large dataset (Stuart A. and Charlie, 2019), and anonymous data was used to pinpoint Muslim cab drivers (Lorenzo, 2015). The attack surface of the exploitation of personal data is not fully discovered. Regularly, new threats appear and are maliciously exploited for influential campaigns and interfering in political election [^3] https://en.wikipedia.org/wiki/Troll_farm, or discovering sensitive information [^4] https://www.nytimes.com/2018/01/29/world/middleeast/strava-heat-map.html.
@@
With this increasing exposure of privacy risks, understanding how young users share their data, and the extent to which they are aware of security and privacy risks, are important properties to assess these risks and to develop effective Privacy and Transparency Enhancing Technologies matching current users’ expectations to increase adoption. Although user perceptions of technology (Beckwith, 2003) and the privacy paradox [^5] The privacy paradox refers to self-reported concerns about privacy appear to be in contradiction with often careless online behaviors. (Barth et al., 2019; Kang and Jung, 2021) have received a lot of attention, users’ self-reported behaviors in mobility contexts associated with smartphones, and the impact of an awareness demonstration platform have been less studied in the academic literature. To fill this gap, this work addresses two research questions:
@@
  • RQ1: What are the perceptions and the understanding of young users’ privacy and its protection in a mobility context?
@@
  • RQ2: What is the impact of a demonstrator for the visualization of location traces and associated privacy risks on these perceptions and understanding?
@@ 2.1. Privacy and location
The privacy issues raised by location data gained a lot of traction in the last decade (Primault et al., 2019). In particular, user location traces extracted from various data have been shown to be highly unique (De Montjoye et al., 2013; Zang and Bolot, 2011). This high uniqueness may act as a digital fingerprint and lead to the re-identification of users if their traces are associated with external knowledge. This uniqueness does not only concern location traces but characterizes all traces generated by a human  (Andreas et al., 2016; Eckersley, 2010; Rebekah and Rachel, 2016), and can generate a risk of re-identification. Several cases of re-identification have been documented, for instance the re-identification of individuals from web search queries (Barbaro and Zeller, 2006), taxi logs (Zhang and Wang, 2016), or the notorious case of Governor William Weld using medical information (Barth-Jones, 2012).
@@
The uniqueness and risk of re-identification is not the unique threat related to location traces. Recent works have demonstrated that location is a very rich contextual information and leads to a strong inferential potential in terms of information that can be learned about individuals (Boutet and Gambs, 2019). For instance, location traces can reveal Points Of Interest (POI) of users such as their home and workplaces (Gambs, Killijian, and del Prado Cortez, 2010), their race and gender (Zhong et al., 2015), their social network (Sharad and Danezis, 2014), it can be used to predict their location patterns (Sadilek and Krumm, 2012), to link accounts of the same user across different datasets (Riederer et al., 2016a), and infer even more sensitive information such as their religion or personality traits [^6] Note that the processing of such sensitive information is prohibited by Article 9(1) of the General Data Protection Regulation (GDPR) on the processing of special categories of personal data..
@@
Due to the large adoption of mobile devices, the location data of users is extensively collected and shared (Almuhimedi et al., 2015; Englehardt and Narayanan, 2016). The uncontrolled usage of this information can have an important impact on users such as unfair price discrimination (Mikians et al., 2012). The increasing use of connected devices collecting personal data and the lack of transparency on how the data is actually exploited raise privacy concerns. Only a handful of tools have been proposed to improve user awareness about the potential risk of revealing their location. For instance, (Riederer et al., 2016b) propose a tool to inspect the potential of location data, while (Boutet and Gambs, 2019) show the impact of protection mechanisms of the inference capabilities using a demonstration platform.
2.2. Perceptions of privacy and privacy controls
@@
Giving users the benefits of location services on their mobile devices while preserving their privacy is an ongoing challenge, as evidenced by mobile OSs iterating on the user interfaces for location notices and control every few years. For example, iOS allows users to select whether apps get fine- or coarse-grained location data [^7] https://support.apple.com/guide/iphone/control-the-location-information-you-share-iph3dd5f9be/ios, and will present pop-ups when apps continually access location services in the background. Similarly, Android apps now have to request background location access separately from general location usage [^8] https://developer.android.com/about/versions/11/privacy/location, and the OS will automatically revoke unused permissions from apps [^9] https://developer.android.com/about/versions/11/privacy/permissions. Even though mobile operating systems regularly improve user interfaces, the opaque privacy controls of location services still face criticism [^10] https://www.attorneygeneral.gov/taking-action/attorney-general-josh-shapiro-announces-391-million-settlement-with-google-over-location-tracking-practices/. (Balash et al., 2022) explore the users’ perceptions regarding access to Google accounts by mobile applications and formulate design recommendations to improve the current third-party management tools offered by Google, such as tracking recent access, automatically revoking access due to app disuse, and providing permission controls. (Wijesekera et al., 2015) analyze authorization preferences in different usage contexts and suggest determining the situations in which users would like to be confronted with security decisions.
@@
Users’ perceptions of the risks and benefits of technologies can determine their willingness to adopt them (Poikela and Kaiser, 2016). More specifically, people are more likely to accept potentially invasive technology if they think its benefits will outweigh its potential risks (Beckwith, 2003). Due to the massive adoption of location-enabled mobile applications, this fact suggests that users’ perception of this trade-off is more in favor of the benefits than the potential risks. This attitude however depends on the perception of the said privacy risks. Studies have shown that users of mobile phones are often unaware of the data collected by apps running on their devices, and that a majority of users restrict some of their permissions (Almuhimedi et al., 2015) following a better awareness of data collection. Other studies explored the privacy paradox (Barth et al., 2019; Kang and Jung, 2021) where self-reported concerns about privacy appear to be in contradiction with often careless online behaviors. However, another study focuses on university community (Gamarra et al., 2019) and show that this population of users does not have a genuine concern regarding the privacy of their geolocation data. Note that this paper did not study location data in mobility contexts associated with smartphones as we do.
@@
Analyzing behavior and understanding users’ perceptions are also important notions to grasp in order to further design Privacy and Transparency Enhancing Technologies (PETs and TETs). For instance, (Kaushik et al., 2021) conducted an online survey to understand people’s perspectives on solely automated decision-making. They then formulate recommendations on how to design such systems. In a similar line of work, (Islami, Fischer-Hübner, and Papadimitratos, 2022) conducted in-depth semi-structured interviews with 17 Swedish drivers to analyse their privacy perceptions and preferences for intelligent transportation systems, then to provide recommendations for suitable predefined privacy options. Several other works (Zufferey et al., 2023; Velykoivanenko et al., 2021; Mink et al., 2022) studied the perceptions of privacy and utility of users related to fitness-trackers and demonstrated a high potential for data minimization (i.e., reducing the volume of data sent to service provider). (Debatin et al., 2009), in turn, investigated Facebook users’ awareness of privacy issues and perceived benefits and risks of utilizing Facebook and recommended better privacy protection, higher transparency and more education about the risks of posting personal information to reduce risky behavior. (Bielova et al., 2024) analysed the behaviors of websites’ visitors through a study of the impact of dark patterns on consent decisions. (Veys et al., 2021) explored whether current data downloads (such as Google Takeout) actually achieve the transparency goals embodied by the right of access. Most participants indicated that current offerings need improvement to be useful, emphasizing the need for better filtration, visualization, and summarizing to help them hone in on key information. However, none of these related work specifically targets users’ behavior and perception about data-sharing of location data. Although (Martin and Nissenbaum, 2019) clearly addressed users’ perceptions of location data, it is to be noted that it does so from an information science standpoint, closer to sociology and not from a UX/usability one. It nonetheless offers a relevant account for the interested reader.
@@
Related to location data, (Farke et al., 2021) specifically analysed the user perceptions and reactions to Google’s My Activity and how this web history dashboard increases or decreases end-users’ concerns and benefits regarding data collection. Their results show that participants were surprised by the volume and detail of the collected data, but most of them were significantly more likely to be both 1) less concerned about data collection and 2) to view data collection more beneficially. However, this dashboard does not present any risks such as possible sensitive inferences associated with location traces or places visited. By also presenting the risks, our study shows that users are more concerned about privacy issues.
3. Methodology
@@
In a nutshell, the first questionnaire (hereinafter “behavior questionnaire”) aimed at assessing their self-reported practices, perceptions, and pre-conceived ideas about privacy. Following this first questionnaire, participants were then invited to use a demonstrator (denoted “demonstration” in what follows) which analyses location traces and reports several inferences as well as the impact of a protection mechanism using differential privacy (i.e., geo-indistiguishability (André s et al., 2013)). The second questionnaire (hereinafter “perception questionnaire”) addressed their assessment of the risks.
3.1. Recruitment
@@ 3.2. Ethical considerations
Second, we elicited answers from participants, which is also personal data. We took as many precautions as needed to guarantee 1) informed consent (see Section A.1 for the consent form, and as noted in Section 3.1 they had the possibility to decline without any penalty), and 2) security of data storage and processing (authors only communicated using end-to-end encrypted tools, and the survey was conducted on an EU-based online tool stamped GDPR-compliant). One participant refused to answer the questionnaires.
@@
3.3. Design of the Questionnaires
We shortly describe in this section the organization of the two questionnaires. The interested reader will find their full content in Appendix A. Each section is succinctly described in what follows:
@@ 3.3. Design of the Questionnaires
3.4. Risks analysis - demonstration
After the first behavior questionnaire, participants had to explore a demonstration platform informing them about privacy risks linked to location traces through example data. We use the (boutethal-02421828) demonstration platform. Participants were asked to use the platform to explore these location traces and to inspect the information that can be inferred. The demonstration typically offers to visualize one’s traces per day (see Figure 1), which can raise awareness about how easy it is to identify Points of Interests (POIs) – such as the home, the workplace or any other visited places during the day –, and presents metadata (e.g., address, category of the place, attendance statistics).
Figure 1. Screenshot of the demonstration’s interface. Clicking on a POI yields attendance statistics.
The demonstration also offers a defense mechanism based on Differential Privacy (Dwork, 2006) in the form of a slider, which controls the level of noise injected into the location data (see Figure 2). This feature is presented as a means to stimulate interest in privacy protection, and to raise awareness of the degradation of data quality when we increase protection with greater noise (to showcase the utility and privacy trade-off).
Figure 2. Illustration of the defense mechanism. Here the red lines show the initial traces, and the black lines represent the traces after the application of the noise.
Finally, the demonstration highlights the risks raised by analyzing location traces and inferring gender, big five personality traits, home, and workplace (see Figure 3). Given the large number of inferences of sensitive information from location data reported in the literature (e.g., people encountered, religious affiliation, sexual orientation, or health status (Gambs, Killijian, and del Prado Cortez, 2010)), only the presentation of these inferences fits fairly well into a representative risk of targeting (i.e., based on predictions computed on data collected and potentially exchanged between different parties).
Figure 3. Inference performed by the platform. The left-hand side shows inference from raw data while the right-hand side presents the reduced risks after the application of the defense mechanism (i.e., geo-indistinguishability (André s et al., 2013)). The noisy version hardly yields the gender, smooths the personality traits, makes the daily and weekly attendance statistics unusable, and infers inaccurate POIs.
3.5. Data analysis
We performed a thematic analysis on the free text answered in the study (see Tables 1 and 2 in Appendix B) to analyse qualitative data. We took inspiration from the steps described in (Braun and Clarke, 2006):
  1. (1)